A magazine reported that 33% of Canadians bike to work. If 300 Canadians are selected at random, find the probability that exactly 111 bike to work.
Answer #1

Solution :

Given that ,

p = 0.33

1 - p = 1 -0.33= 0.67

n = 300

x = 111

Using binomial probability formula ,

P(X = x) = ((n! / (n - x)!) * px * (1 - p)n - x

P(X = 111) = ((300 ! / (300 - 111)!) * 0.33111 * 0.67300 - 111

= 0.0164

Probability =0.0164